What is the molar solubility of Fe(OH)2 when buffered at pH of 9.00? [Fe(OH)2 : Ksp = 7.9 x 10^-16]
Answer #1

Fe(OH)2 ---------> Fe2+ + 2OH-

pH = 9

pOH = 14- pH = 14 - 9

pOH = 5

[OH] = 10-pOH = 10-5

-------------------------------------

Ksp = [Fe2+].[OH-]2

[Fe2+] = (Ksp) / ([OH-]2)

Therefore, Molar solubility is given by:

[Fe2+] = (7.9 x 10-16) / (10-5)2

[Fe2+] = 7.9 x 10-6 M -------------------(ANSWER)